OpenStudy (anonymous):

Can you explaing me how: tan^-1(2/2)=pi/4 How became tan^-1(2/2) in pi/4. originary exercise complex number z^9=2+2j

OpenStudy (mathmath333):

can u tell what is the orignal question

OpenStudy (anonymous):

Ok one momente a make a photo.

OpenStudy (mathmath333):

\[\tan^{-1}(\dfrac{2}{2})\\=\tan^{-1}(1)=45^{\circ}=\dfrac{\pi}{4}\]

OpenStudy (anonymous):

ok and how 45 grade in pi/4?

OpenStudy (anonymous):

its on the unit circle

OpenStudy (mathmath333):

w

OpenStudy (mathmath333):

suppose this is right angled triangle and here it is also iscoceles triangle so the two sides opposite to \(45^{\circ}\) are equal now we know that \(\tan \theta =\dfrac{opposite~ side}{adjacent~side}\) so \(\tan 45^{\circ} =\dfrac{a}{a}\) \(\tan 45^{\circ} =1\\\) hence \(45^{\circ}=\tan ^{-1} 1\)

OpenStudy (mathmath333):

and \(45^{\circ}\) in radian is \((\dfrac{\pi}{4})\)

OpenStudy (anonymous):

can I make it without calculator?

OpenStudy (anonymous):

or onliy wiht calclator?

OpenStudy (mathmath333):

which thing ?

OpenStudy (anonymous):

a sorry radians

OpenStudy (mathmath333):

well the idea is that \(\pi\) radian is equal to \(180^{\circ}\) to \(~~~~~\dfrac{\pi}{4}\\ =\dfrac{180}{4}\\ =\dfrac{90}{2}\\ =45^{\circ}\)

OpenStudy (mathmath333):

or conversly \(\pi\) radian is equal to \(180^{\circ}\) \(~~~~~45^{\circ}\\ =45\times \dfrac{\pi}{180}\\ =\dfrac{\pi}{4}\\ \)

OpenStudy (anonymous):

ok only one thing about it

OpenStudy (anonymous):

than I divide 45pi/180 by 45 ?

OpenStudy (anonymous):

a ok I understand. also Pi must be alone without number?

OpenStudy (mathmath333):

\(\pi\) is not alone it has some value like \(3.14..\) but we usually write angles in it in the degrees \(\Huge \theta ^{\circ}\) symbol \(\huge \pi\)

OpenStudy (anonymous):

yes ok that is so

OpenStudy (anonymous):

but I mean the number before pi

OpenStudy (anonymous):

by every exercise liek this I must divide from Pi. the pi must stay alone wihtout other numbers?

OpenStudy (mathmath333):

for example if u have to to write \(360^{\circ}\) in radian then it can be written as \(\Large 360^{\circ}\\ \Large =360\times \dfrac{\pi}{180}\\ \Large =2\pi\\\)

OpenStudy (anonymous):

a ok I understand thank you very much what arguments I must study for this exerscise? I understand all but I will study it for sin and cos

OpenStudy (mathmath333):

here is a good link for u http://www.mathsisfun.com/geometry/unit-circle.html

OpenStudy (mathmath333):

at first trignometry is difficult,but once u practice , it will be as easy as \(1+1=2\)

OpenStudy (anonymous):

Ok will do but first I must study relevant things for test. thanky you

OpenStudy (mathmath333):

welcome